# A Novel Formula for Solving Quadratic Equations over Binary Extension Fields

### v1

@@ Abstract
Solving quadratic equations over finite fields is a fundamental task in algebraic coding theory and serves as a key subroutine for computing the roots of cubic and quartic polynomials. For the reduced quadratic polynomial $x^{2}+x+c\in\mathbb{F}_{2^{m}}[x]$, existing formula-based methods rely on heavy exponentiation or case distinctions on $m$ (odd/even or powers of two), which limits uniformity and efficiency. This paper presents a unified, formula-based solution for all positive integers $m$ that uses only exclusive-OR operations (XORs). The approach leverages a Reed-Muller matrix characterization of evaluations and reduces the problem to solving a binary linear system. The total cost is at most $m^{2}-2m+1$ XORs, and under parallelism, the latency is $\lceil\log_{2}m\rceil$ XORs, making the method attractive for low-power, low-latency applications.
@@ I Introduction
Finding the roots of polynomials over finite fields is a fundamental task in algebraic coding theory. In Bose-Chaudhuri-Hocquenghem (BCH) and Reed-Solomon (RS) codes [lin2001error], for example, the roots of the error-locator polynomial directly determine error positions. A widely used approach is the Chien search [chien1964], which evaluates the polynomial at every field element and identifies points at which the evaluation vanishes. Although its hardware-friendly structure enables low latency and broad adoption, exhaustive evaluation incurs substantial computational redundancy and, consequently, avoidable power consumption.
It is classical that polynomials of degree at most four over the complex field admit formula-based solutions in terms of their coefficients. Analogously, certain low-degree polynomials over finite fields also permit direct, non-exhaustive root computation, with quadratic equations playing a central role. Particularly, roots of cubic and quartic polynomials can be reduced to solving associated quadratic instances (see, e.g., [chen1982formulas]). Consequently, the unified, formula-based solution for quadratic polynomials developed here can be leveraged to obtain roots of cubic and quartic polynomials over $\mathbb{F}_{2^{m}}$. For brevity, this paper focuses on the quadratic case and omits higher-degree derivations.
@@
If $f(x)$ has roots in $\mathbb{F}_{2^{m}}$, then it must be two distinct roots, i.e., $x_{0},x_{1}$, due to $x_{0}+x_{1}=1$. It should be noted that the formula solution of $f(x)\in\mathbb{F}_{2^{m}}[x]$ was first proposed in [chen1982formulas]. However, the proposed method in [chen1982formulas] is not unified, as $m$ must be split into odd and even cases for separate treatment. Moreover, the even case needs to be further partitioned by $m\bmod 4$, complicating implementation. A different approach was later proposed in [cherly1998solving], which depends on finding an element whose trace function value is one. Furthermore, [walker1999new] proposed a formula to solve the reduced polynomial, but the proposed one applies only when $m$ is a power of two. Notably, all of the above works rely heavily on exponentiation in $\mathbb{F}_{2^{m}}$, in addition to field additions, which hampers efficiency and uniform implementation across all $m$. As a comparison, this paper proposes a unified, formula-based solution for quadratic equations over $\mathbb{F}_{2^{m}}$, where $m$ is an arbitrary positive integer. Importantly, the proposed method requires only a small number of XOR operations, avoiding exponentiation. The comparison details between the proposed method and previous methods are shown in TABLE II.
@@ II Solving Quadratic Equations
Since $\ell$ is a power of two, one can know from formulas (13) and (15) in [yu2023reed] that $\begin{pmatrix}w_{0}^{\ell}&w_{1}^{\ell}&w_{2}^{\ell}&\cdots&w_{2^{m}-1}^{\ell}\end{pmatrix}=E_{\ell}\cdot R_{m}$, where $E_{\ell}$ is the row vector whose $2^{j}$-th element is $\alpha^{j\ell}$ and the other elements are zero. The above formula gives (3). This completes the proof. ∎
@@ II-A Proposed Method
which, if solvable over $\mathbb{F}_{2^{m}}$, has two distinct roots $x_{0}$ and $x_{1}$ satisfying $x_{0}+x_{1}=1$. The following lemma characterizes solvability and identifies the roots via a binary linear system.
@@
where $R_{m}$ and $R_{m}(2^{j})$ are defined in Lemma 1, and $R_{m}(0)$ denotes the first row of $R_{m}$. The above formula is similar to the encoding formula of first-order Reed-Muller (RM) codes (please refer to [dumer2006soft, reeves2023reed] for details).
@@
where the $\ell$-th row, $0\leq\ell<m$, is exactly the codeword of the first-order RM code corresponding to the message vector
@@
From [dumer2006soft, reeves2023reed], the message polynomial corresponding to (10) is as follow: (each $i_{j},0\leq j<m,$ represents an independent variable)
@@
Notably, the matrix $B$ in Lemma 2 is independent of $f(x)$, so it can be pre-calculated, and one can record the process of transforming $B$ into its reduced row echelon form. Precisely, let $P$ be the row transformation matrix that transforms $B$ into reduced row echelon form. Since the first column of $B$ is a zero column and $B$ has the rank of m1m-1 (which is due to the equation of (6) having at most two distinct roots), the reduced row echelon form of $B$ must be
@@
Given the matrix pair $(P,I_{0})$ for $\mathbb{F}_{2^{m}}$, which is shown in (14) and where the $\ell$-th row of $I_{0}$ is an all-zero row, then $f(x)$ has roots in $\mathbb{F}_{2^{m}}$ if and only if the $\ell$-th element of the following result is also zero:
@@
has a sulotion on $i_{1},i_{2},\ldots,i_{m-1}$. Since the $\ell$-th rows of $\mathbf{s}$ and $I_{0}$ are all zeros, $(i_{1},i_{2},...,i_{m-1})$ can easily be obtained to maintain the above identity. Note that $I_{0}$ is the $m\times(m-1)$ matrix obtained by inserting one all-zero row to an $(m-1)\times(m-1)$ identity matrix. Thus, $f(x)$ must have roots. Conversely, if $f(x)$ has roots and $f(w_{i})=0$, then (6) in Lemma 2 holds, which gives (16). Since the $\ell$-th row of $I_{0}$ is an all-zero row, the $\ell$-th element of $\mathbf{s}$ must be zero. This completes the proof.
Theorem 2.
Given the matrix pair $(P,I_{0})$ for $\mathbb{F}_{2^{m}}$, which is shown in (14), then $f(w_{i})=0$ and $f(w_{i}+1)=0$, where $i=\sum_{j=1}^{m-1}i_{j}\cdot 2^{j}$, satisfies
Proof.
II-B Example
It can be transformed into the reduced row echelon form
through the following process (row indexing starts from zero):
  1. Step 1)
    add $1$-th row to $2$-th row, add $3$-th row to $6$-th row.
  2. Step 2)
    add $2$-th row to $4$-th row, add $5$-th row to $3$-th rows.
  3. Step 3)
    add $4$-th row to $1$-th and $2$-th rows.
  4. Step 4)
    add $6$-th row to $1$-th, 22-th, 33-th, and $5$-th rows.
Obviously, the matrix $I_{0}$ is the matrix obtained by adding an all-zero row on top of the $6\times 6$ identity matrix. In addition, the above process leads to the row transformation matrix $P$ in the matrix pair $(P,I_{0})$ being
According to the formula solution in Theorem 2,
@@ 4. Step 4)
[chien1964] arbitrary None (using exhaustive search). Field Add, Mul: $2^{m+1}$, $2^{m}$
[chen1982formulas] odd $x_{0}=\sum_{j\in J}c^{2^{j}}=\sum_{i\in I}c^{2^{i}}$, where Field Exponentiation: (m1)/2(m-1)/2
@@
[cherly1998solving] arbitrary $x_{0}=\sum_{j=1}^{m-1}c^{2^{j}}\cdot\left(\sum_{\ell=0}^{j-1}u^{2^{\ell}}\right)$,where $u\in\mathbb{F}_{2^{m}}$ and Field Exponentiation: m1m-1
@@
[walker1999new] power of two There are formulas corresponding to distinct $m$. Field Mul: 22 ($m=2$)
@@ III Complexity Analysis
This section analyzes the computational complexity of the above method. It should be noted that the matrix pair $(P,I_{0})$ in Theorems 1 and 2 is determined only based on the finite field $\mathbb{F}_{2^{m}}$. The computational complexity of checking whether the reduced quadratic polynomial $f(x)$ has solutions (i.e., Theorem 1) is as follows:
@@
In the following, the method proposed in this paper is compared with other methods in [chien1964, chen1982formulas, cherly1998solving, walker1999new]. The third column of TABLE II provides the formula expression for one root $x_{0}$ (clearly, the other root is $x_{1}=x_{0}+1$), and the fourth column shows the number of operations required for each method. Note that for the case where $m$ is an even number in [chen1982formulas], only the operations for calculating $Tr_{4}$(c) were counted, although the remaining computational costs are substantial. In [cherly1998solving], the field element $u$ can be found in advance, and its related results can be pre-calculated. Furthermore, the method in [walker1999new] needs to transform $f(x)$ into the form of $x^{2}+t_{0}x+t_{1}$, where $t_{0}\in\mathbb{F}_{2^{m}}\setminus\mathbb{F}_{2^{m/2}}$, $t_{1}\in\mathbb{F}_{2^{m}}$, and then apply the corresponding formula to solve. This generates a large number of operations that can be pre-calculated. All pre-calculated operation quantities are not included in TABLE II. Due to the fact that field addition is the simplest and most efficient implementation among all operations (a field addition over $\mathbb{F}_{2^{m}}$ is equivalent to $m$ XORs), one can see that the proposed method not only has the uniform form for arbitrary $m$, but also operates most efficiently, requiring only $m^{2}-2m+1$ XORs.

### v2

@@ Abstract
Solving quadratic equations over finite fields is a fundamental task in algebraic coding theory and serves as a key subroutine for computing the roots of cubic and quartic polynomials. Notably, any quadratic polynomial over binary extension fields can be transformed into the reduced form $x^{2}+x+c\in\mathbb{F}_{2^{m}}[x]$, for which existing formula-based methods rely on heavy exponentiation or case distinctions on $m$ (odd/even or powers of two), limiting uniformity and efficiency. This paper presents a unified, formula-based solution for all positive integers $m$ that uses only exclusive-OR operations (XORs). The approach leverages a Reed-Muller matrix characterization of evaluations and transforms the problem into computing a binary matrix-vector multiplication. The total cost is at most $m^{2}-2m+1$ XORs, and under parallelism, the latency is $\lceil\log_{2}m\rceil$ XORs, making the method attractive for low-power, low-latency applications.
@@ I Introduction
Finding the roots of polynomials over finite fields is a fundamental task in algebraic coding theory. In Bose-Chaudhuri-Hocquenghem (BCH) and Reed-Solomon (RS) codes [5], for example, the roots of the error-locator polynomial directly determine error positions. A widely used approach is the Chien search [3], which evaluates the polynomial at every field element and identifies points at which the evaluation vanishes. Although its hardware-friendly structure enables low latency and broad adoption, exhaustive evaluation incurs substantial computational redundancy and, consequently, avoidable power consumption.
It is classical that polynomials of degree at most four over the complex field admit formula-based solutions in terms of their coefficients. Analogously, certain low-degree polynomials over finite fields also permit direct, non-exhaustive root computation, with quadratic equations playing a central role. Particularly, roots of cubic and quartic polynomials can be reduced to solving associated quadratic instances (see, e.g., [1]). Consequently, the unified, formula-based solution for quadratic polynomials developed here can be leveraged to obtain roots of cubic and quartic polynomials over $\mathbb{F}_{2^{m}}$. For brevity, this paper focuses on the quadratic case and omits higher-degree derivations.
@@
If $f(x)$ has roots in $\mathbb{F}_{2^{m}}$, then it must be two distinct roots, i.e., $x_{0},x_{1}$, due to $x_{0}+x_{1}=1$. It should be noted that the formula solution of $f(x)\in\mathbb{F}_{2^{m}}[x]$ was first proposed in [1]. However, the proposed method in [1] is not unified, as $m$ must be split into odd and even cases for separate treatment. Moreover, the even case needs to be further partitioned by $m\bmod 4$, complicating implementation. A different approach was later proposed in [2], which depends on finding an element whose trace function value is one. Furthermore, [7] proposed a formula to solve the reduced polynomial, but the proposed one applies only when $m$ is a power of two. Notably, all of the above works rely heavily on exponentiation in $\mathbb{F}_{2^{m}}$, in addition to field additions, which hampers efficiency and uniform implementation across all $m$. As a comparison, this paper proposes a unified, formula-based solution for quadratic equations over $\mathbb{F}_{2^{m}}$, where $m$ is an arbitrary positive integer. Importantly, the proposed method requires only a small number of XOR operations, avoiding exponentiation. The comparison details between the proposed method and previous methods are shown in TABLE II.
@@ II Solving Quadratic Equations
Since $\ell$ is a power of two, one can know from formulas (13) and (15) in [8] that $\begin{pmatrix}w_{0}^{\ell}&w_{1}^{\ell}&w_{2}^{\ell}&\cdots&w_{2^{m}-1}^{\ell}\end{pmatrix}=E_{\ell}\cdot R_{m}$, where $E_{\ell}$ is the row vector whose $2^{j}$-th element is $\alpha^{j\ell}$ and the other elements are zero. The above formula gives (3). This completes the proof. ∎
@@ II-A Proposed Method
which, if solvable over $\mathbb{F}_{2^{m}}$, has two distinct roots $x_{0}$ and $x_{1}$ satisfying $x_{0}+x_{1}=1$. One can obtain the following lemma.
@@
where $R_{m}$ and $R_{m}(2^{j})$ are defined in Lemma 1, and $R_{m}(0)$ denotes the first row of $R_{m}$. The above formula is similar to the encoding formula of first-order Reed-Muller (RM) codes (please refer to [4, 6] for details).
@@
where the $\ell$-th row, $\forall 0\leq\ell<m$, is exactly the codeword of the first-order RM code corresponding to the message vector
@@
From [4, 6], the message polynomial corresponding to (10) is as follows: (each $i_{j},0\leq j<m,$ represents an independent variable)
@@
Notably, the matrix $B$ in Lemma 2 is independent of $f(x)$, so it can be pre-calculated, and one can record the process of transforming $B$ into its reduced row echelon form. Let $P$ be the row transformation matrix that transforms $B$ into reduced row echelon form. Since the first column of $B$ is a zero column and $B$ has the rank of m1m-1 (which is due to the equation of (6) having at most two distinct roots), the reduced row echelon form of $B$ must be
@@
Given the matrix pair $(P,I_{0})$ for $\mathbb{F}_{2^{m}}$, which is shown in (14) and where the $\ell$-th row of $I_{0}$ is an all-zero row, $f(x)$ has roots in $\mathbb{F}_{2^{m}}$ if and only if the $\ell$-th element of the following result is zero:
@@
has a solution on $i_{1},i_{2},\ldots,i_{m-1}$. Since the $\ell$-th rows of $\mathbf{s}$ and $I_{0}$ are all zeros, $(i_{1},i_{2},...,i_{m-1})$ can easily be obtained. Thus, $f(x)$ must have roots. Conversely, if $f(x)$ has roots and $f(w_{i})=0$, then (6) in Lemma 2 holds, which gives (16). Since the $\ell$-th row of $I_{0}$ is an all-zero row, the $\ell$-th element of $\mathbf{s}$ must be zero. This completes the proof.
@@
[3] arbitrary None (using exhaustive search). Field Add, Mul: $2^{m+1}$, $2^{m}$
[1] odd $x_{0}=\sum_{j\in J}c^{2^{j}}=\sum_{i\in I}c^{2^{i}}$, where Field Exponentiation: (m1)/2(m-1)/2
@@
[2] arbitrary $x_{0}=\sum_{j=1}^{m-1}c^{2^{j}}\cdot\left(\sum_{\ell=0}^{j-1}u^{2^{\ell}}\right)$,where $u\in\mathbb{F}_{2^{m}}$ and Field Exponentiation: m1m-1
@@
[7] power of two There are formulas corresponding to distinct $m$. Field Mul: 22 ($m=2$)
@@
Theorem 2.
Given the matrix pair $(P,I_{0})$ for $\mathbb{F}_{2^{m}}$, which is shown in (14), then $f(w_{i})=0$ and $f(w_{i}+1)=0$, where $i=\sum_{j=1}^{m-1}i_{j}\cdot 2^{j}$ satisfies
Proof.
II-B Example
It can be transformed into the reduced row echelon form
through some elementary row transformations. Obviously, the matrix $I_{0}$ is the matrix obtained by adding an all-zero row on top of the $6\times 6$ identity matrix. In addition, the above process leads to the row transformation matrix $P$ in the matrix pair $(P,I_{0})$ being
According to the formula solution in Theorem 2,
@@ III Complexity Analysis
This section analyzes the computational complexity of the above method. The computational complexity of checking whether the reduced quadratic polynomial $f(x)$ has solutions (i.e., Theorem 1) is as follows:
@@
In the following, the method proposed in this paper is compared with other methods in [3, 1, 2, 7]. The third column of TABLE II provides the formula expression for one root $x_{0}$ (clearly, the other root is $x_{1}=x_{0}+1$), and the fourth column shows the number of operations required for each method. Note that for the case where $m$ is an even number in [1], only the operations for calculating $Tr_{4}$(c) were counted, although the remaining computational costs are substantial. In [2], the field element $u$ can be found in advance, and its related results can be pre-calculated. Furthermore, the method in [7] needs to transform $f(x)$ into the form of $x^{2}+t_{0}x+t_{1}$, where $t_{0}\in\mathbb{F}_{2^{m}}\setminus\mathbb{F}_{2^{m/2}}$, $t_{1}\in\mathbb{F}_{2^{m}}$, and then apply the corresponding formula to solve. This generates a large number of operations that can be pre-calculated. All pre-calculated operation quantities are not included in TABLE II. Due to the fact that field addition is the simplest and most efficient implementation among all operations (a field addition over $\mathbb{F}_{2^{m}}$ is equivalent to $m$ XORs), one can see that the proposed method not only has the uniform form for arbitrary $m$, but also operates most efficiently, requiring only $m^{2}-2m+1$ XORs.